An 108.0 kg boy and a 61.1 kg girl, both wearing skates face each other at rest on a skating rink. The boy and girl push each other, sending her eastward with a speed of 3.09 m/s. Determine the final velocity of the boy. Neglect friction.


1
Expert's answer
2022-04-15T16:40:14-0400

"m_bv_b+m_gv_g=0,"

"v_b=-\frac{m_g}{m_b}v_g=-\frac{61.1}{108}\cdot 3.09=-1.75\space m/s."

Answer: 1.75 m/s westward.